OpenStudy (anonymous):

Need help with 10th grade geometry word problems,,,

OpenStudy (anonymous):

A rectangle is 5 times as long as it is wide. The peremiter is 110cm. Find the dimensions of the rectangle. Round to nearest thenth if necessary

OpenStudy (saifoo.khan):

The formula to find the perimeter: 2(L+B) Length: 5x Width: x Now insert the values in the formula.

OpenStudy (saifoo.khan):

Perimeter = 2(L+B) 110 = 2(5x + x) Solve for x.

OpenStudy (anonymous):

we have 4 possible answers... none look like that... son forgot book in locker - we have 3 word problems

OpenStudy (anonymous):

The sides of a triangle are in the ratio 3:4:5 What is the length of each side if the perimeter of the triangle is 42

OpenStudy (saifoo.khan):

That's not the answer, you have to solve it further..

OpenStudy (anonymous):

ok ill give that to him to try and figure out more....

OpenStudy (anonymous):

can u help with other 2 problems? The sides of a triangle are in the ratio 3:4:5 What is the length of each side if the perimeter of the triangle is 42

OpenStudy (anonymous):

please

OpenStudy (saifoo.khan):

The sides of a triangle are in the ratio 3:4:5 What is the length of each side if the perimeter of the triangle is 42 3x + 4x + 5x = 42 Solve for x. Then substitute the value of x in all 3 terms.

OpenStudy (anonymous):

OK last one for today... Two cars leave Denver at the same time and travel in opposite direction. one car travels 5mi/h faster that the other car. The care are 575 mi apart in 5 hours. How fast is each car traveling

OpenStudy (saifoo.khan):

@yummydum

OpenStudy (saifoo.khan):

Oh man! Duhh me.

OpenStudy (saifoo.khan):